Rotten Tomato: A bag contains 9 tomatoes, of which 1 is rotten. A sample of three tomatoes is selected at random. What is the probability that the sample contains the rotten tomato?

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

If you put the 1st one back each time, it's 1/9 + 1/9 + 1/9 = 1/3.
If you don't, it's 1/9 + 1/8 + 1/7 = 191/504.
------------------
I corrected this, 1st thought it said 10 tomatoes.

Answer by ewatrrr(24785)


Hi
9 tomatoes, of which 1 is rotten. three tomatoes is selected at random
P(1 rotten) = 1 - P(no rotten) = 1 - 8C3/9C3) = 1 - .6677 = .333